OpenStudy (anonymous):

What is the derivative of 5(3x)^.5

OpenStudy (anonymous):

\[7.5(3x)^{-1/2}\]

OpenStudy (anonymous):

\[5*3X=15\]

OpenStudy (anonymous):

please show me how you got 15 step by step

OpenStudy (anonymous):

@nish933

OpenStudy (anonymous):

i mean 15X

OpenStudy (anonymous):

then

OpenStudy (anonymous):

0.5*15X^-1/2

OpenStudy (anonymous):

7.5X\[7,5X^-\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

did u get it @anthonykanow

OpenStudy (cwrw238):

you use the chain rule for (3x)^0.5 = 3* 0.5(3x)^-0.5 then multiply by 5 = 7.5(3x)^-0.5

OpenStudy (anonymous):

No I did not. @nish933 please use mathprint. \[5(3x)^{1/2}\] \[2.5(3x)^{-1/2}\] \[(3)(2.5)(3x)^{-1/2}=7.5(3x)^{^{-1/2}}\]

OpenStudy (cwrw238):

yep thats right